   There are several ways how to relax the property that a graph possesses a
hamiltonian path/cycle. It is usually done if we are not able to show that the
graph is hamiltonian, or if we want to measure how close the graph is to being
 Hamiltonian.\ \In this talk we will present problems and results of the above
     type. In particular, we will discuss the famous Middle Layer Problem.